1. At a price of $200, a cellphone company manufactures 300,000 units. At a price of $150, the company produces 200,000 phones. What is the price elasticity of supply?
2. Do customers who visit convenience stores at 3 a.m. have a price elasticity of demand that is more or less elastic than those who visit at 3 p.m.? Please explain.
1
Expert's answer
2019-12-23T11:24:56-0500

1. Price, p1 =$200, for Q1 300000 units

Price, p2 =$150, for Q2 200000 units.

Price elasticity of supply,"PEs= \frac{\Delta Q}{\Delta P}"

"ΔQ=(Q2−Q1)/\frac{Q2+Q1}{2}" =-100000/250000 =-0.4

"ΔP=(P2−P1)/\frac{P2+P1}{2}" = -50/175= -0.3

"PEs=\frac{−0.4}{−0.3}" =1.33

2. Customers who visit convenience stores at 3a.m. likely have a price elasticity of demand that is less elastic than those who visit at 3 p.m. This is largely because the 3 a.m. costumers have fewer substitutes for available convenience stores, which means that they cannot change their quantity demanded as much as the 3 p.m. people can in response to changes in prices